If the perimeter of a rectangle is 238 and the length is 75 more than the width, what is the length and the width
Answer by lefty4ever26(59)

L=75+W
Perimeter means add up all the sides. L+L+W+W. Replace the L's with (75+W)
(75+W)+(75+W)+W+W=238
150+4W=238
4W=88
W=22 SO if the Width is 22 then the Length is 22+75=97.